Find the solution to the system by the substitution methond.
y=4x-4
y=5x
a.(1,0)
b.(-4,-20)
c.(4,0)
d.(-3,-15)
(Im confused !!!please explain to me how get the answer)!

Answer by ankor@dixie-net.com(22740)

Find the solution to the system by the substitution methond.
y=4x-4
y=5x
:
Where you see y in the first equation substitute 5x (from the 2nd equation)
5x = 4x - 4
subtract 4x from both sides
5x - 4x = -4
x = -4
:
Now you know that x = -4, substitute -4 for x in the 2nd equation
y = 5(-4)
y = -20
:
The x,y solution then is -4, -20
:
:
A good idea to check your solutions in the 1st equation, substitute for both x and y, see if it is true
-20 = 4(-4) - 4
-20 = -16 -4 which is true.